the width of a rectangle is two less than the length.If the perimeter of the rectangle is 28 meters,find the dimension of the rectangle?
Answer by l-mayfield(2)

The equation for P=2L+2W We know that W is 2 less than L => W=L-2 we can now plug that into the equation to solve. 28=2L + 2(L-2) => 28 = 2L + 2L -4 you can rearrange the equation to read 2L + 2L -4 = 28. Add 4 to both sides and combine like terms => 4L = 32 now divide both sides by 4 to get L=8 and W =6
Test: 2(8)+2(6) = 16 + 12 =28